# A Generalization of von Neumann’s Reduction from the Assignment Problem to Zero-Sum Games

### v2

@@ 5.2 Production Planning Game
We now provide a more general, natural interpretation of the reduced game for any linear program with positive $b$ and $c$. Such linear programs can be viewed as planning optimal production. A central authority decides how much of each of $n$ goods to produce, where $x_{j}$ is the amount of good $j$ produced. The profit from producing one unit of good $j$ is $c_{j}$. The production of each good consumes certain amounts of $m$ resources representing, e.g., raw materials and electricity, or, more broadly, labor hours and machine capacity. Producing a unit of the $j$th good requires $a_{ij}$ units of the $i$th resource. Finally, we have total availability of $b_{i}$ units for the $i$th resource. The total usage of resources is not allowed to exceed their availability, which is exactly captured by the constraints $Ax\leq b$. Furthermore, it is reasonable to assume positive profits from producing a good as well as a positive availability of each resource. Hence, the production planning scenario is captured by the linear program $(P)$ for positive $b$ and $c$. In many applications, we also have nonnegative entries in $A$. However, negative entries in $A$ allow for modeling further scenarios, such as producing a resource, e.g., electricity, as a byproduct in a production process.
@@
The reduced game now has the following interpretation: the column player mimics the production planner while the row player mimics a resource provider. A pure strategy for the column player is to pick a good while the row player picks a resource. While the production planner wants to gain high production profits, the resource player wants to sell them a scarce resource. Selecting the $j$th good and $i$th resource leads to a payment by the column player to the row player of $\frac{a_{ij}}{b_{i}c_{j}}$. This quantity is the resource share needed to produce a unit of the $j$th good, $\frac{a_{ij}}{b_{i}}$, divided by the profit $c_{j}$ for producing one unit. Together, this is the required resource share per unit profit. If the selected good leads to a small profit or needs a high amount of the resource, i.e., the resource is scarce for this good, then the payment is high. Conversely, needing a small share of the resource or making a lot of profit leads to a small payment, or even a negative payment if the resource is produced as a byproduct indicated by a negative $a_{ij}$. Hence, it makes intuitive sense that the payoff depends on the reciprocals of both the resource availability and the price of the good.
5.3 Playing Linear Programs
We want to broaden the insights of the discussed applications to give an even more general interpretation of our correspondence.
For comparison, we first recall the original construction by Dantzig (1951) as defined in Section 3 given by the game matrix in Equation 1. The skew-symmetric matrix of this game leads to symmetric roles for both players. They both specify strategies of the form $(p,q,t)$, where $q$ corresponds to a solution of the primal linear program $(P)$ and $p$ corresponds to a solution of the dual linear program $(D)$.
By contrast, consider now our construction for positive $b$ and $c$ as detailed in Theorem 4.3. There, the column and row players have different roles corresponding to the primal and dual linear programs, respectively. Moreover, each fixed strategy gives a guarantee to a player, i.e., a minimum reward for the row player against adversarial play of the column player or a maximum payment by the column player against an adversarial row player. These guarantees are obtained by solving $(\mbox{{Row}})$ and $(\mbox{{Col}})$ when fixing strategies $p$ and $q$, respectively. When players play optimal strategies, they get their optimal guarantee, which is exactly the maximin or minimax value. Importantly, such guarantees are obtained for any strategy, not only when players play optimally and they carry information about the underlying LP.
First, if the column player can play a (not necessarily optimal) strategy $q$ that guarantees her negative utility, i.e., for this fixed $q$, the minimum $v$ of $(\mbox{{Col}})$ is negative, then this $q$ gives rise to an unboundedness certificate as in Part 1-(b) of the proof of Theorem 4.3. Interestingly, the unboundedness can not only be certified by an optimal strategy but by any strategy of the column player yielding nonpositive payoff.
Conversely, any strategy that guarantees a positive payoff leads to a feasible solution to the corresponding linear program. Specifically, assume that $\bar{q}$ and $\bar{p}$ are strategies for the column and row players such that $(\mbox{{Col}})$ and $(\mbox{{Row}})$ have positive optimal values $\bar{v}$ and $\bar{u}$ when fixing $\bar{q}$ and $\bar{p}$, respectively. Then, as in Part 1-(b) of the proof of Theorem 4.3, scaling these strategies as $\frac{1}{\bar{v}}\bar{q}$ and $\frac{1}{\bar{u}}\bar{p}$ leads to feasible solutions of $(P^{\prime})$ and $(D^{\prime})$, respectively. Their objective values with respect to $(P^{\prime})$ and $(D^{\prime})$ are $\frac{1}{\bar{v}}$ and $\frac{1}{\bar{u}}$. Once again, this works even for nonoptimal strategies of the reduced games as long as they guarantee a positive payment. However, importantly, such a strategy exists for the row player only if $(D^{\prime})$ admits a feasible solution, and, therefore, $(P^{\prime})$ is not unbounded.
All these considerations give a broader interpretation of the game play as playing a linear program: In the first case, the column player plays an infeasibility certificate. In the second case, the row player can play a feasible strategy, implying that any strategy of the column player has to lead to a positive value of $(\mbox{{Col}})$. Hence, in this case, both players play a feasible solution. This solution might, however, lead to a suboptimal objective value of the LP. In the special case where both play maximin strategies, their play corresponds to playing optimal solutions of the LP, which we can extract alongside the optimal value of the LP.
Notably, as in the case of the assignment problem, the direct correspondence, where solutions only differ by scaling, is obtained between the linear program pair $(P^{\prime},D^{\prime})$ and the zero-sum game based on the constraint matrix of this game. In other words, linear programs of the form $(P^{\prime},D^{\prime})$ are linear programs of games. Von Neumann’s result can therefore be interpreted as the realization that the assignment problem can be written in game form. And more generally, Proposition 4.1 shows that, whenever $b>\mathbf{0}$ and $c>\mathbf{0}$, any linear program pair of the form $(P,D)$ can be written in game form, as well.
6 Relationship with Dantzig’s Reduction
@@
Recall our discussion in Section 1 that Dantzig’s original reduction does not work for the full generality of linear programs. In this section, we explore whether the restrictions that we impose on linear programs suffice to make his reduction work. First, we show that under the assumption $b>\mathbf{0}$ and $c>\mathbf{0}$, his reduction works. Let $(p^{*},q^{*},t^{*})$ be the minimax strategy of the row player for Dantzig’s game as presented in Equation 1. Then, it holds that
$(i)$ $Aq^{*}-bt^{*}\leq\mathbf{0},$
$(ii)$ $-A^{\intercal}p^{*}+ct^{*}\leq\mathbf{0},$
$(iii)$ $b^{\intercal}p^{*}-c^{\intercal}q^{*}\leq\mathbf{0},$
$(iv)$ $\mathbf{1}^{\intercal}p^{*}+\mathbf{1}^{\intercal}q^{*}+t^{*}=1,$
$(v)$ $p^{*}\geq\mathbf{0},\;q^{*}\geq\mathbf{0},\;t^{*}\geq 0,$
$(vi)$ $t^{*}(b^{\intercal}p^{*}-c^{\intercal}q^{*})=0.$
@@
The following statements are true:
@@
    If $t^{*}>0$, then $\left(\frac{1}{t^{*}}q^{*},\frac{1}{t^{*}}p^{*}\right)$ is a primal-dual optimal pair for $(P,D)$.
@@
    We first consider the case $t^{*}>0$. Conditions ($i$), (iiii), and ($v$) imply that $\frac{1}{t^{*}}q^{*}$ and $\frac{1}{t^{*}}p^{*}$ are feasible for $(P)$ and $(D)$, respectively. In addition, by ($vi$), $\frac{1}{t^{*}}\left(b^{\intercal}p^{*}-c^{\intercal}q^{*}\right)=0$, which completes the necessary and sufficient conditions for the optimality of $\frac{1}{t^{*}}q^{*}$ and $\frac{1}{t^{*}}p^{*}$.
@@
    Now consider the case $t^{*}=0$. Assume for contradiction that $q^{*}=0$. Then, by ($iv$) and ($v$), it holds that $0\neq p^{*}\geq 0$. Hence, since $b>\mathbf{0}$, it follows that $b^{\intercal}p^{*}>0$, contradicting (iiiiii). Thus, $q^{*}\neq 0$ which implies (as $c>\mathbf{0}$) that $c^{\intercal}q^{*}>0$. Moreover, by ($i$) and ($v$), it holds that $Aq^{*}\leq 0$ and $q^{*}\geq 0$. Together, $q^{*}$ is an unboundedness certificate for (P)(P).∎
Notably, the second part of Proposition 6.1 has the precondition $t^{*}=0$ whereas the theorem by Dantzig (1951) had the precondition $b^{\intercal}p^{*}-c^{\intercal}q^{*}<0$. Recall that the “hole” in Dantzig’s proof was due to the fact that it could happen that both $t^{*}=0$ and $b^{\intercal}p^{*}-c^{\intercal}q^{*}=0$. This raises the question whether Proposition 6.1 works because this case never occurs or whether it works because this case can be dealt with successfully (unlike in Dantzig’s reduction). Remarkably, the latter is the case, as our next example shows. Regardless of conditions on $p^{*}$ and $q^{*}$, we obtain an unboundedness certificate whenever $t^{*}=0$.
@@
The matrix of Dantzig’s reduced zero-sum game is given by the matrix
$\begin{pmatrix}0&0&-1&\frac{1}{2}&-1\\ 0&0&1&-\frac{1}{2}&-1\\ 1&-1&0&0&1\\ -\frac{1}{2}&\frac{1}{2}&0&0&1\\ 1&1&-1&-1&0\\ \end{pmatrix}\text{.}$
@@
Consider any constraint matrix $A\geq\mathbf{0}$ and consider the linear program pair given by $A$, $b=\mathbf{0}$, and $c=\mathbf{0}$. Dantzig’s reduced game as in Equation 1 for this case is
@@
Arguably, the case of $b=c=\mathbf{0}$ in the last example seems somewhat trivial. However, we want to stress that Dantzig’s correspondence can only fail for somewhat trivial cases, namely cases covered by our preprocessing in Algorithm 1. Indeed, instead of using our reduction, we can also use Dantzig’s reduction in Algorithm 1 for tackling LPs with $A\geq\mathbf{0}$. This approach is sound due to Proposition 6.1.
7 Computational Considerations
Up to this point, we have omitted the computational properties of reductions and the number spaces to which they are applicable. In this section, we address these aspects.
7.1 Reasonable Reductions
Typically, in complexity theory, meaningful reductions are required to be polynomial, that is, the effort of the reduction, as well as the effort to recover the solution, is required to be polynomial in the size of the problem. We shall assume for now that the data of the linear programs is rational and discuss implications of our reduction for real input data subsequently. To make the reduction meaningful in the framework of complexity theory and given that any LP problem with rational data is known to be solvable in time polynomial in the size of the problem, we follow Adler (2013) and require reductions to be strongly polynomial, that is, they have to use a polynomial number of arithmetic operations and, if applied to rational input data, polynomial space (Tardos, 1986; Grötschel et al., 1993). In our case, the input data of a linear program depends on its dimensions $m$ and $n$ and the encoding of numbers, hence polynomial means with respect to $m$, $n$, and the logarithm of numbers such as the entries of the constraint matrix. Note that strongly polynomial reductions still permit the occurrence of exponentially large numbers with respect to $m$ and $n$, and hence all the previously existing reductions by Adler (2013), Brooks and Reny (2023), and von Stengel (2023) are strongly polynomial as well. In the context of linear programming, strongly polynomial reductions make a lot of sense since it is a long-standing open problem to decide whether linear programs can be solved in strongly polynomial time (see, e.g., Tardos, 1986; Grötschel et al., 1993, Section 6.6).
Notably, our as well as previous reductions from linear programs to zero-sum games are even log-space reductions, i.e., they only use logarithmic space except for outputting the polynomially large reduced instance (Arora and Barak, 2009). Note that even a log-space reduction can produce exponentially large numbers with respect to input parameters, such as the large numbers in the reductions by Adler (2013), Brooks and Reny (2023), and von Stengel (2023). Their reductions run in log-space because these numbers do not have to be computed with arithmetic operations but can directly be inferred from the input parameters and outputted without computation. [^9] For example, recall the number $m(\sqrt{m}\theta)^{m}+1$ from Footnote 4. In binary representation, this number only uses a polynomial number of bits with respect to the input size. Instead of being computed it can be inferred directly from the input data and simply outputted as a matrix entry of the reduced game. An interesting feature of previous reductions being log-space is that it implies that computing maximin strategies in zero-sum games is P-complete because solving linear programs is P-complete (Dobkin et al., 1979). Apparently, this observation was missed in previous work.
7.2 Reductions for Real Numbers
In the previous subsection, we assumed rational data for the discussion of reasonable reductions. However, in the beginning of Section 2, the problems are presented in terms of real numbers. This presentation is common whenever the mathematics of linear programming problems are discussed. The reason for this is that, as recognized from the early days of linear programming and game theory, all major mathematical theorems (e.g., the LP-Duality Theorem or the Minimax Theorem) are valid for real numbers. [^10] In fact, they are valid for any ordered field (see, e.g., Gale, 1960)). In addition, most related algorithms, such as Fourier-Motzkin elimination or the simplex algorithm, prior to the polynomial-time algorithm for solving LPs by Khachiyan (1979) were developed and presented in terms of real numbers.
Indeed, both the reductions by Dantzig (1951) and von Neumann (1953) only use elementary computations and are applicable for real numbers. However, the corrected extensions of Dantzig’s reduction of Adler (2013), Brooks and Reny (2023), and von Stengel (2023) are only applicable for rational LPs because the constructions require a big number depending on the data. It is currently not known whether this number can be calculated in strongly polynomial time for real-valued LPs. [^11] By contrast, the Turing-reduction by Adler (2013) is applicable for real input but it requires solving up to $n$ LPs.
Not being applicable to real numbers feels like a missing feature of these reductions. LPs and zero-sum games in real numbers are generally relevant for complexity analysis of computational models with real numbers (see, e.g., Blum et al., 1989). In particular, the problem of finding a strongly polynomial algorithm for real-valued LPs is listed as one of the most prominent mathematical challenges for the 21st century (Smale, 1998, Problem 9).
Interestingly, our reduction shares the applicability for real numbers with the original reduction of von Neumann. It thus provides a somewhat more complete equivalence between LPs and zero-sum games, albeit only for positive $b$ and $c$ or nonnegative $A$. Note that only reductions from linear programs to zero-sum games that are applicable to real numbers can be used to translate strongly polynomial algorithms (for real input) for finding maximin strategies for zero-sum games to strongly polynomial algorithms to solve linear programs. In other words, using our technique, any strongly polynomial algorithm for computing maximin strategies for zero-sum games would immediately give rise to a strongly polynomial algorithm for solving linear programs with positive $b$ and $c$ or nonnegative $A$.
8 Discussion
Clearly, the question of generalizing von Neumann’s result (von Neumann, 1953) to an arbitrary primal-dual pair of LPs is an important one. Our work explores an extension based on rescaling, the crucial idea that helps define the payoffs in the hide-and-seek game. Even though this seems only possible if $b>\mathbf{0}$ and $c>\mathbf{0}$, we managed to reduce the case $A\geq\mathbf{0}$ to this case as well. Going further may well require a new approach.
Our paper is first and foremost an attempt at expounding a purely technical result; yet, its historical overtones are hard to miss. In that vein, we would like to end with some observations which may interest students and observers of the history of computing and computability.
A lot has been said about von Neumann’s extraordinary intellectual prowess and his foundational influence on the creation of the digital computer, in particular, his 1945 report on what is known today as the von Neumann architecture (see, e.g., Halmos, 1973; Bhattacharya, 2021). Our study of his masterful work (von Neumann, 1953) has led us to believe that he had an intuitive understanding of the notion of efficient computability, i.e., polynomial-time solvability.
With a view to solving the assignment problem efficiently, von Neumann (1953) gave the following procedure: Reduce the assignment problem to a zero-sum game and approximately solve the latter using methods from Brown and von Neumann (1951), involving solving a sequence of differential equations.
Over the centuries, other mathematicians had also found algorithms that were later seen to be polynomial-time algorithms, e.g., Euclid’s algorithm for computing the greatest common divisor of two numbers. However, von Neumann’s work is qualitatively different. He explicitly reasoned about the time complexity of the method he had proposed, as is evident from reading the first page of the paper (von Neumann, 1953). The page starts by observing that an optimal solution to the assignment problem can be found via an exhaustive search, i.e., by considering all $n!$ permutation matrices, and ends by stating that his procedure gives a “significant approximate solution” in “a moderate power of nn” steps. The page ends with the following striking sentence, which is clearly alluding to the qualitative difference between polynomial and exponential time.
It turns out that this number is a moderate power of $n$, i.e., considerably smaller than the “obvious” estimate $n!$ mentioned earlier.
Additionally:
  1. 1.
    Von Neumann (1953) reduced the assignment problem to the problem of finding a Nash equilibrium in zero-sum games.
  2. 2.
    Today we know that both these problems are in P: the first due to the work of Kuhn (1955) and the second because the problem is equivalent to linear programming, which is in P due to Khachiyan (1979); moreover, the reduction is polynomial time. Recall that the class P was defined more than a decade after von Neumann’s work (Edmonds, 1965).
Hence, von Neumann gave a strongly polynomial-time reduction between two problems in P — over a decade before P was even defined!

### v3

@@ 4.1 Production Planning Games
LPs with positive $b$ and $c$ can be viewed as planning optimal production: A central authority decides how much of each of $n$ goods to produce, where $x_{j}$ is the amount of good $j$ produced. The profit from producing one unit of good $j$ is $c_{j}$. The production of each good consumes certain amounts of $m$ resources representing, e.g., raw materials and electricity, or even labor hours and machine capacity. Producing a unit of the $j$th good requires $a_{ij}$ units of the $i$th resource. Finally, we have total availability of $b_{i}$ units for the $i$th resource. The total usage of resources is not allowed to exceed their availability, which is exactly captured by the constraints $Ax\leq b$. Furthermore, it is reasonable to assume positive profits from producing a good as well as a positive availability of each resource. Hence, the production planning scenario is captured by the linear program $(P)$ for positive $b$ and $c$. In many applications, we also have nonnegative entries in $A$. However, negative entries in $A$ allow for modeling further scenarios, such as producing a resource, e.g., electricity, as a byproduct in a production process.
@@
The reduced game now has the following interpretation: the column player mimics the production planner while the row player mimics a resource provider. A pure strategy for the column player is to pick a good while the row player picks a resource. While the production planner wants to gain high production profits, the resource player wants to sell them a scarce resource. Selecting the $j$th good and $i$th resource leads to a payment by the column player to the row player of $\frac{a_{ij}}{b_{i}c_{j}}$. This quantity is the resource share needed to produce a unit of the $j$th good, that is, $\frac{a_{ij}}{b_{i}}$, divided by the profit $c_{j}$ for producing one unit. Together, this is the required resource share per unit profit. If the selected good leads to a small profit or needs a high amount of the resource, i.e., the resource is scarce for this good, then the payment is high. Conversely, needing a small share of the resource or making a lot of profit leads to a small payment, or even a negative payment if the resource is produced as a byproduct indicated by a negative $a_{ij}$. Hence, it makes intuitive sense that the payoff depends on the reciprocals of both the resource availability and the price of the good.
This perspective also offers a novel view on the hide-and-seek game. In this context, the production planner generates edges. For instance, the planner could be an event manager organizing tennis matches. A match between players $i$ and $j$ yields a surplus $\phi_{ij}$. The resource provider supplies the players, who correspond to the vertices in the assignment problem. By the rules of the hide-and-seek game, if a player participates in a desired match, then a compensation of $1/\phi_{ij}$ must be paid. Assuming that the surplus from a match accrues linearly over its duration, with a total surplus of $\phi_{ij}$, this compensation corresponds to the fraction of the match that yields one unit of profit.
4.2 Computational Considerations
The key idea of the papers repairing Dantzig’s original reduction is to modify his reduced game: Adler (2013) and von Stengel (2023) add an additional row while Brooks and Reny (2023) transform the game with elementary column operations. These reductions have in common the fact that the constraint matrix or vectors in the reduced games need large numbers, dependent on the input numbers and exponential in $m$ or nn. [^4] As an example, in the reduction by Adler (2013), the additional entry of the constraint vector is estimated as a number that is at least $m(\sqrt{m}\theta)^{m}+1$ where $\theta$ is the absolute value of the largest number contained in $A$ and $b$. Nonetheless, since the large numbers can be computed with a polynomial number of arithmetic operations and have a bit size polynomial in the input bit size, the reductions are strongly polynomial. [^5] For a formal definition of strongly polynomial time, we refer to Grötschel et al. (1993). In addition, Adler (2013) also provides a Turing reduction based on sequential variable elimination that avoids large numbers. Our reduction is a strongly polynomial reduction without large numbers.
Within the different realm of computational models based on Turing machines, all of the aforementioned reductions are log-space reductions, i.e., they only use logarithmic space except for outputting the polynomially large reduced instance (Arora and Barak, 2009). In fact, the occurring large numbers can be computed bit-wise in logarithmic space. This has the interesting consequence that computing maximin strategies in zero-sum games is P-complete because solving linear programs is P-complete (Dobkin et al., 1979). Apparently, this observation was missed in previous work.
4.3 Relationship with Dantzig’s Reduction
The reduction in the seminal paper by Dantzig (1951) is based on the following skew-symmetric game matrix, given the input data of the primal LP (P)(P). [^6] Dantzig (1951) attributes the idea for this matrix to Tucker.
Denoting a maximin strategy as $(p^{*},q^{*},t^{*})$, the key result of the paper by Dantzig (1951) states:
  1. 1.
[…]
  2. 2.
@@ 4.3 Relationship with Dantzig’s Reduction
In this section, we explore whether the restrictions that we impose on linear programs suffice to make his reduction work. First, we show that under the assumption $b>\mathbf{0}$ and $c>\mathbf{0}$, his reduction works. Let $(p^{*},q^{*},t^{*})$ be a minimax strategy of the row player. Then, it holds that
$\begin{array}[]{lll}(i)\ Aq^{*}-bt^{*}\leq\mathbf{0},&(ii)\ -A^{\intercal}p^{*}+ct^{*}\leq\mathbf{0},&(iii)\ b^{\intercal}p^{*}-c^{\intercal}q^{*}\leq\mathbf{0},\\ (iv)\ \mathbf{1}^{\intercal}p^{*}+\mathbf{1}^{\intercal}q^{*}+t^{*}=1,&(v)\ p^{*}\geq\mathbf{0},\;q^{*}\geq\mathbf{0},\;t^{*}\geq 0,&(vi)\ t^{*}(b^{\intercal}p^{*}-c^{\intercal}q^{*})=0.\end{array}$
@@
Assume that $b>\mathbf{0}$ and $c>\mathbf{0}$ and let $(p^{*},q^{*},t^{*})$ be a minimax strategy of the row player in Dantzig’s game given by Equation 5. Then the following statements are true:
@@
    If $t^{*}>0$, then $\left(\frac{1}{t^{*}}q^{*},\frac{1}{t^{*}}p^{*}\right)$ is an optimal primal-dual pair for $(P,D)$.
@@
    We assume $t^{*}>0$. Conditions ($i$), (iiii), and ($v$) imply that $\frac{1}{t^{*}}q^{*}$ and $\frac{1}{t^{*}}p^{*}$ are feasible for $(P)$ and $(D)$, respectively. In addition, by ($vi$), $\frac{1}{t^{*}}\left(b^{\intercal}p^{*}-c^{\intercal}q^{*}\right)=0$, which completes the necessary and sufficient conditions for the optimality of $\frac{1}{t^{*}}q^{*}$ and $\frac{1}{t^{*}}p^{*}$.
@@
    Now assume $t^{*}=0$. Assume for contradiction that $q^{*}=\mathbf{0}$. Then, by ($iv$) and ($v$), it holds that $0\neq p^{*}\geq\mathbf{0}$. Hence, since $b>\mathbf{0}$, it follows that $b^{\intercal}p^{*}>0$, contradicting (iiiiii). Thus, $q^{*}\neq 0$ which implies (as $c>\mathbf{0}$) that $c^{\intercal}q^{*}>0$. Moreover, by ($i$) and ($v$), it holds that $Aq^{*}\leq\mathbf{0}$ and $q^{*}\geq\mathbf{0}$. Together, $q^{*}$ is an unboundedness certificate for (P)(P).∎
Notably, the second part of Proposition 4.1 has the precondition $t^{*}=0$ whereas the theorem by Dantzig (1951) had the precondition $b^{\intercal}p^{*}-c^{\intercal}q^{*}<0$. Recall that the issue with Dantzig’s proof was that it could happen that both $t^{*}=0$ and $b^{\intercal}p^{*}-c^{\intercal}q^{*}=0$. This raises the question whether Proposition 4.1 works because this case never occurs or whether it works because this case can be dealt with successfully (unlike in Dantzig’s reduction). Remarkably, the latter is the case, as our next example shows. Regardless of conditions on $p^{*}$ and $q^{*}$, we obtain an unboundedness certificate whenever $t^{*}=0$.
@@
Consider any constraint matrix $A\geq\mathbf{0}$ and consider the linear program pair given by $A$ as well as $b=\mathbf{0}$ and $c=\mathbf{0}$. Dantzig’s reduced game as in Equation 5 for this case is
@@
Arguably, the case of $b=\mathbf{0}$ and $c=\mathbf{0}$ in the last example may seem unsatisfying. However, we want to stress that Dantzig’s correspondence can only fail for somewhat trivial cases, namely cases covered by our preprocessing in Algorithm 1. Indeed, instead of using our reduction, we can also use Dantzig’s reduction in algorithm 1 of Algorithm 1 for tackling LPs with $A\geq\mathbf{0}$. This approach is sound due to Proposition 4.1.